# Generalized Entropy Calibration for Inference with Partially Observed Data: A Unified Framework

### v1

@@ 3.1 Dual formulation
The constrained optimization (3.2)–(3.4) is solved efficiently through its dual. Introduce Lagrange multipliers $\lambda_{1}\in\mathbb{R}^{q}$ and $\lambda_{2}\in\mathbb{R}$ for constraints (3.3)–(3.4), and write $b_{i}=b(O_{i})$, $\hat{g}_{i}=g(\hat{\pi}_{i}^{-1})$, and $s_{i}=(b_{i}^{\top},\hat{g}_{i})^{\top}$. The Lagrangian is
$Q(\omega,\lambda)$ $=-\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i})+\lambda_{1}^{\top}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b_{i}-\sum_{i=1}^{N}b_{i}\right)+\lambda_{2}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,\hat{g}_{i}-\sum_{i=1}^{N}\hat{g}_{i}\right),$ (3.5)
where $\lambda=\left(\lambda_{1}^{\top},\lambda_{2}\right)^{\top}$. Maximizing $Q(\omega,\lambda)$ with respect to $\omega_{i}$ yields the closed-form optimal weights
$\omega_{i}^{*}(\lambda)=g^{-1}(\lambda_{1}^{\top}b_{i}+\lambda_{2}\hat{g}_{i})=g^{-1}(\lambda^{\top}s_{i}),$ (3.6)
where $\lambda=(\lambda_{1}^{\top},\lambda_{2})^{\top}$ and $g^{-1}$ is strictly increasing because $G$ is strictly convex. The function $g(\cdot)$ is called the calibration link function and is closely related to the canonical link function in the generalized linear models. The calibration link operates on the weight parameter $\omega_{i}$ rather than a conditional mean $\mu_{i}$, but the algebraic structure is identical. Substituting (3.6) back to (3.5) gives the dual objective
$\rho_{G}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,F(\lambda^{\top}s_{i})-\frac{1}{N}\sum_{i=1}^{N}\lambda^{\top}s_{i},$ (3.7)
where $F(\nu)=-G\{g^{-1}(\nu)\}+g^{-1}(\nu)\,\nu$ is the convex conjugate of $G$. The optimal dual solution
$\hat{\lambda}=\arg\min_{\lambda}\rho_{G}(\lambda)$
@@
Differentiating (3.7) confirms that $\hat{\lambda}$ satisfies the calibration equations (3.3)–(3.4):
@@ 4 Large-Sample Properties
We establish the asymptotic properties of the proposed estimator $\hat{\theta}_{\omega}$ defined in (3.1). Throughout, the estimator solves
@@
$\nabla\ell(\varphi)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\biggl(\frac{\delta_{i}}{\pi_{i}}-1\biggr)h_{i}(\varphi)\;=\;0,$ (4.1)
$\nabla\hat{\rho}_{G}(\lambda)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\Biggl\{\delta_{i}\,\omega_{i}^{*}(\lambda,\hat{\varphi})\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}-\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}\Biggr\}\;=\;0,$ (4.2)
@@
The regularity conditions are collected in the Supplementary Material (SM), Section A. Assumptions 1–2 ensure the consistency of $\hat{\varphi}$ and $\hat{\lambda}$, established in Lemma S.1 of the SM. Assumptions 3–4 impose smoothness and non-degeneracy of the limiting dual and propensity-score objectives; in particular, non-singularity of hessian $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}$, where $\lambda^{*}=\arg\min_{\lambda}\rho_{G}(\lambda)$ and $\rho_{G}(\lambda)$ is defined in (3.7), requires that the calibration covariates $s_{i}=(b_{i}^{\top},g_{i})^{\top}$ are not collinear among the respondents, which is the natural identifiability condition for entropy calibration. Assumptions 5–7 are standard conditions for joint M-estimation of the system (4.1)–(4.2); they hold whenever the three sub-problems are individually well-posed and their interaction is smooth.
@@ 4.1 Linearization
The following theorem provides a first-order linear expansion of $\hat{U}^{\omega}(\theta)$ that underlies all subsequent asymptotic results.
@@ 4.2 Consistency under the correct PS model
We first show that the calibration weights reduce to IPW weights when the PS model is correct, and derive the resulting asymptotic distribution.
@@
$V_{1}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\left\{U(\theta_{0};Z)-\gamma^{*}S(O;\varphi_{0})-\kappa^{*}h(O;\varphi_{0})\right\}^{\otimes 2}\right],$
@@
$\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\varphi}\biggl\{\gamma^{*}s_{i}(\varphi)+\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\varphi)\bigl[U(\theta;\,z_{i})-\gamma^{*}s_{i}(\varphi)\bigr]+\Bigl(1-\frac{\delta_{i}}{\pi(O_{i};\varphi)}\Bigr)\kappa\,h_{i}(\varphi)\biggr\}=0.$ (4.3)
@@ 4.3 Consistency under the correct OR model
If the OR model is correctly specified, i.e., $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$, and $b^{*}(O)=\mathbb{E}\{U(\theta;Z)\mid O\}$ is used in (3.3), then $\kappa^{*}=0$.
@@
$\bar{V}_{1}\;=\;\mathrm{Var}\bigl\{\mathbb{E}[U(\theta_{0};Z)\mid O]\bigr\}\;+\;\mathbb{E}\!\Big[\delta\,\bigl\{\omega^{*}(O;\lambda^{*},\varphi^{*})\bigr\}^{2}\,\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\Big].$
@@ 4.4 Double robustness, variance dominance, and local efficiency
Corollaries 1 and 2 together establish that $\hat{\theta}_{\omega}$ is doubly robust: it is consistent for $\theta_{0}$ if either the PS or the OR model is correctly specified. We now show that, under a correct PS model, the proposed estimator is never less efficient than the classical AIPW estimator and is strictly more efficient whenever the OR model is misspecified and the debiasing covariate carries additional information.
@@
$V_{3}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\left\{U(\theta_{0};Z)-b(O)\right\}^{\otimes 2}\right]$
@@
The variance reduction is most pronounced when the OR model is substantially misspecified, so that $b(O)$ is a poor approximation to $\mathbb{E}\{U(\theta_{0};Z)\mid O\}$, but the PS model provides useful information about the missingness mechanism through the debiasing covariate $g(\pi^{-1}(O;\varphi_{0}))$ and the score function $h(O;\varphi_{0})$. In such settings, the entropy calibration framework effectively enriches the augmentation space beyond what AIPW uses, yielding strictly smaller asymptotic variance without requiring any additional modeling effort.
@@
$V_{1}^{*}\;=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\right],$
@@
5 Computational Details
This section describes the numerical implementation of $\hat{\theta}_{\omega}$. The dual formulation in Section 3.1 reduces the calibration step to an unconstrained convex minimization over $\lambda\in\mathbb{R}^{q+1}$. We now address two additional computational issues: (i) solving the nested optimization when the calibration function $b(O_{i})$ depends on $\theta$, and (ii) constructing $b(O_{i})$ via cross-fitted machine-learning predictions.
@@ 5.1 Profile optimization
$\hat{\theta}_{\omega}=\arg\max_{\theta}\min_{\lambda}\;\rho_{G}(\lambda,\theta).$ (5.1)
Equivalently, defining the profile objective $L(\theta)=\rho_{G}(\hat{\lambda}(\theta),\theta)$ where $\hat{\lambda}(\theta)=\arg\min_{\lambda}\rho_{G}(\lambda,\theta)$, we have $\hat{\theta}_{\omega}=\arg\max_{\theta}L(\theta)$.
We solve (5.1) using a two-loop procedure. For a fixed $\theta$, the inner loop minimizes $\rho_{G}(\lambda,\theta)$ over $\lambda$ using Newton’s method. The outer loop updates $\theta$ via a Newton or BFGS step based on the gradient of $L(\theta)$. By the envelope theorem (Danskin, 2012), $\nabla_{\theta}L(\theta)=\nabla_{\theta}\rho_{G}(\hat{\lambda}(\theta),\theta)$, since $\nabla_{\lambda}\rho_{G}=0$ at the inner optimum. The procedure is summarized in Algorithm 1.
Algorithm 1 Two-Loop Profile Optimization for $\hat{\theta}_{\omega}$
@@
5:  Outer loop: Update $\theta^{(k+1)}\leftarrow\theta^{(k)}-[\nabla^{2}_{\theta}L(\theta^{(k)})]^{-1}\,\nabla_{\theta}L(\theta^{(k)})$
@@ 5.2 Calibration via cross-fitting
In practice, the conditional expectation $b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$ is unknown and must be estimated. Following the idea of Angelopoulos et al. (2023), we approximate it by plugging in a prediction $\hat{M}_{i}$ of the missing component $M_{i}$: $b^{*}(\theta;O_{i})\approx U(\theta;\,O_{i},\hat{M}_{i})$, where $\hat{M}_{i}$ is obtained from a flexible machine-learning model fitted to the observed data. To avoid the overfitting bias that arises from using the same data to train the prediction rule and to estimate $\theta$ (Chernozhukov et al., 2018), we adopt $K$-fold cross-fitting. Let $\mathcal{I}=\{1,\ldots,N\}$ and $\mathcal{S}=\{i\in\mathcal{I}\mathrel{\mathop{\ordinarycolon}}\delta_{i}=1\}$. We randomly partition $\mathcal{I}$ into $K$ disjoint folds $\{\mathcal{I}^{(k)}\}_{k=1}^{K}$. For each fold $k$, we fit a prediction model $\hat{m}^{(-k)}(\cdot)$ using training data $\mathcal{S}^{(-k)}=\mathcal{S}\setminus(\mathcal{S}\cap\mathcal{I}^{(k)})$ and compute out-of-fold predictions $\hat{M}_{i}^{(-k)}=\hat{m}^{(-k)}(O_{i})$ for $i\in\mathcal{I}^{(k)}$. Using these cross-fitted predictions, we solve the calibration problem:
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (5.2)
$\text{subject to}\quad\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}\delta_{i}\,\omega_{i}\,U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right)=\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right),$ (5.3)
@@
Note that the prediction model $\hat{m}^{(-k)}$ is trained only on the complete cases in $\mathcal{S}^{(-k)}$, but the out-of-fold predictions $\hat{M}_{i}^{(-k)}$ are computed for all units $i\in\mathcal{I}^{(k)}$, including those with $\delta_{i}=0$. This is essential: the balancing constraint (5.3) involves the full sample, so calibration function values are needed for every unit regardless of whether $M_{i}$ is observed.
@@
The asymptotic results in Section 4 require the cross-fitted predictor $\hat{m}^{(-k)}$ to satisfy a mean-squared-error condition of the form $\mathbb{E}\|\hat{M}_{i}^{(-k)}-\mathbb{E}(M_{i}\mid O_{i})\|^{2}=o(1)$, i.e., the prediction error must vanish in probability. This condition is mild and is satisfied by a broad range of modern machine-learning methods, including random forests, gradient boosting, neural networks, and penalized regression, provided the sample size grows and standard regularity conditions hold. No specific convergence rate (e.g., $N^{-1/4}$) is required for the cross-fitted predictions, because the debiasing constraint ensures that the estimator remains consistent even when the prediction model converges slowly, as long as the PS model is correctly specified. When neither model is correctly specified, faster convergence of the predictor generally improves finite-sample performance, though consistency is no longer guaranteed. See Section C of SM for more details.

### v2

@@ 3.1 Dual formulation
The primal program in (2)–(4) expresses the statistical idea of the method, but the resulting weights are most conveniently characterized through its dual. The dual formulation serves two purposes. First, it yields a closed-form representation of the optimal weights in terms of a low-dimensional multiplier $\lambda$, making clear how the balancing and debiasing constraints jointly determine the final calibration rule. Second, it reduces the constrained optimization over $N$ weights to an unconstrained convex optimization over $\lambda\in\mathbb{R}^{q+1}$, which is the key computational advantage of generalized entropy calibration.
To derive the dual representation, introduce Lagrange multipliers $\lambda_{1}\in\mathbb{R}^{q}$ and $\lambda_{2}\in\mathbb{R}$ for constraints (3)–(4), where $q$ denotes the dimension of $b(O)$, including its intercept term. Write $b_{i}=b(O_{i})$, $\hat{g}_{i}=g(\hat{\pi}_{i}^{-1})$, and $s_{i}=(b_{i}^{\top},\hat{g}_{i})^{\top}$. The Lagrangian is
$Q(\omega,\lambda)$ $=-\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i})+\lambda_{1}^{\top}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b_{i}-\sum_{i=1}^{N}b_{i}\right)$
$\qquad+\lambda_{2}\left(\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,\hat{g}_{i}-\sum_{i=1}^{N}\hat{g}_{i}\right),$ (5)
where $\lambda=(\lambda_{1}^{\top},\lambda_{2})^{\top}$.
Maximizing $Q(\omega,\lambda)$ with respect to $\omega_{i}$ yields the closed-form optimal weights
$\omega_{i}^{*}(\lambda)=g^{-1}(\lambda_{1}^{\top}b_{i}+\lambda_{2}\hat{g}_{i})=g^{-1}(\lambda^{\top}s_{i}),$ (6)
where $g^{-1}$ is strictly increasing because $G$ is strictly convex. The function $g(\cdot)$ is called the calibration link function and is closely related to the canonical link function in the generalized linear models. The calibration link operates on the weight parameter $\omega_{i}$ rather than a conditional mean $\mu_{i}$, but the algebraic structure is identical.
Substituting (6) back to (5) gives the dual objective
$\rho_{G}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,F(\lambda^{\top}s_{i})-\frac{1}{N}\sum_{i=1}^{N}\lambda^{\top}s_{i},$ (7)
where $F(\nu)=-G\{g^{-1}(\nu)\}+g^{-1}(\nu)\,\nu$ is the convex conjugate of $G$. The calibration step is therefore reduced to the convex optimization problem
$\hat{\lambda}=\arg\min_{\lambda}\rho_{G}(\lambda),$
@@
Differentiating (7) confirms that $\hat{\lambda}$ satisfies the calibration equations (3)–(4):
@@ 4 Large-sample properties
This section establishes the large-sample behavior of the proposed estimator $\hat{\theta}_{\omega}$ and clarifies how the calibration construction reproduces the classical doubly robust logic. The analysis proceeds in three steps. We first derive a first-order linearization showing that the estimator admits a calibration-based augmentation representation without requiring either working model to be correct. We then study the two one-model regimes separately: correctness of the propensity-score model and correctness of the outcome-regression model. These results together imply double robustness. Finally, we compare the proposed estimator with the classical AIPW estimator and show that, under a correct propensity-score model, generalized entropy calibration is never less efficient and is strictly more efficient whenever the debiasing covariate contributes information beyond the augmentation space used by AIPW.
Throughout, $\hat{\theta}_{\omega}$ solves
@@
$\nabla\ell(\varphi)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\biggl(\frac{\delta_{i}}{\pi_{i}}-1\biggr)h_{i}(\varphi)\;=\;0,$ (8)
$\nabla\hat{\rho}_{G}(\lambda)$ $\;\equiv\;\frac{1}{N}\sum_{i=1}^{N}\Biggl\{\delta_{i}\,\omega_{i}^{*}(\lambda,\hat{\varphi})\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}-\begin{pmatrix}b_{i}\\ g_{i}(\hat{\varphi})\end{pmatrix}\Biggr\}\;=\;0,$ (9)
@@
The regularity conditions are collected in the Supplementary Material (SM), Section A. Assumptions 1–2 ensure the consistency of $\hat{\varphi}$ and $\hat{\lambda}$, established in Lemma S.1 of the SM. Assumptions 3–4 impose smoothness and nondegeneracy of the limiting dual and propensity-score objectives. In particular, nonsingularity of the Hessian $E\{\nabla^{2}\rho_{G}(\lambda^{*})\}$ requires that the calibration covariates $s_{i}=(b_{i}^{\top},g_{i})^{\top}$ are not collinear among the respondents, which is the natural identifiability condition for entropy calibration. Assumptions 5–7 are standard conditions for joint $M$-estimation of the coupled system defining $\hat{\theta}_{\omega}$, $\hat{\lambda}$, and $\hat{\varphi}$.
@@ 4.1 Linearization
The key technical step is a first-order expansion of $\hat{U}_{\omega}(\theta)$ showing that the proposed estimator behaves like an augmentation estimator in which the augmentation term is generated by calibration. This representation is the basis for all subsequent results and already reveals the role of the calibration covariates $S(O;\phi^{*})$ in separating full-sample information from respondent-only residual variation.
@@ 4.2 Consistency under the correct PS model
We begin with the regime in which the propensity-score model is correctly specified. In this case, the debiasing constraint forces the calibration weights to recover the IPW solution, so the proposed estimator inherits the consistency of inverse-probability weighting while potentially benefiting from additional variance reduction through the calibration structure.
@@
$V_{1}\;$ $=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}\;+\;\mathbb{E}\!\Bigl[\Bigl(\frac{1}{\pi(O;\varphi_{0})}-1\Bigr)$
$\qquad\Bigl\{U(\theta_{0};Z)-\gamma^{*}S(O;\varphi_{0})-\kappa^{*}h(O;\varphi_{0})\Bigr\}^{\otimes 2}\Bigr],$
@@
$\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\varphi}\biggl\{\gamma^{*}s_{i}(\varphi)+\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\varphi)\bigl[U(\theta;\,z_{i})-\gamma^{*}s_{i}(\varphi)\bigr]$
$\qquad+\Bigl(1-\frac{\delta_{i}}{\pi(O_{i};\varphi)}\Bigr)\kappa\,h_{i}(\varphi)\biggr\}=0.$
@@ 4.3 Consistency under the correct OR model
We next consider the complementary regime in which the outcome-regression model is correctly specified. Here the balancing constraint aligns the calibration function with the conditional mean of the estimating function, so the leading bias term is removed even if the propensity-score model is misspecified.
@@
If the OR model is correctly specified, i.e., $\mathbb{E}\{U(\theta_{0};Z)\mid O\}\in\mathrm{span}\{b(O)\}$, and $b^{*}(O)=\mathbb{E}\{U(\theta;Z)\mid O\}$ is used in (3), then $\kappa^{*}=0$.
@@
$\bar{V}_{1}\;$ $=\;\mathrm{Var}\bigl\{\mathbb{E}[U(\theta_{0};Z)\mid O]\bigr\}$
$\qquad\;+\;\mathbb{E}\!\Big[\delta\,\bigl\{\omega^{*}(O;\lambda^{*},\varphi^{*})\bigr\}^{2}\,\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\Big].$
@@ 4.4 Double robustness, variance dominance, and local efficiency
Corollaries 1 and 2 together establish that $\hat{\theta}_{\omega}$ is doubly robust: it is consistent for $\theta_{0}$ whenever either the PS model or the OR model is correctly specified. Double robustness alone, however, does not distinguish the proposed estimator from classical AIPW, which shares this property. The more distinctive question is comparative: once the PS model is correct, does the calibration formulation offer any efficiency advantage over AIPW? The answer is yes, and the reason is geometric.
To see why, recall from Theorem 4.1 that $\hat{\theta}_{\omega}$ behaves asymptotically like an augmentation estimator whose augmentation term is the projection of $U(\theta_{0};Z)$ onto the calibration covariate space
$\mathcal{S}(O;\varphi_{0})=\mathrm{span}\bigl\{b(O),\;g\bigl(\pi^{-1}(O;\varphi_{0})\bigr),\;h(O;\varphi_{0})\bigr\}.$
The classical AIPW estimator, by contrast, projects onto the strictly smaller space $\mathrm{span}\{b(O)\}$ alone. The variance of an augmentation estimator is determined by the residual after projection: a larger projection space yields a smaller residual, and hence a smaller variance. This is the content of Corollary 3 below.
To state the result precisely, define a weighted seminorm on functions of $(O,Z)$ by
$\|f\|_{w}^{2}=\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)f^{\otimes 2}\right],$
where the weight $\pi^{-1}(O;\varphi_{0})-1$ is the excess weight assigned to complete cases under IPW. The variance components of both estimators can be expressed in terms of this seminorm: $V_{3}-V_{1}$ equals the squared $\|\cdot\|_{w}$-distance between the AIPW projection and the GEC projection of $U(\theta_{0};Z)$ onto their respective augmentation spaces. Since $\mathrm{span}\{b(O)\}\subseteq\mathcal{S}(O;\varphi_{0})$, the Pythagorean identity in $\|\cdot\|_{w}$ gives
$\bigl\|U-\Pi_{\{b\}}U\bigr\|_{w}^{2}=\bigl\|U-\Pi_{\mathcal{S}}U\bigr\|_{w}^{2}+\bigl\|\Pi_{\mathcal{S}}U-\Pi_{\{b\}}U\bigr\|_{w}^{2},$ (10)
where $\Pi_{A}$ denotes $\|\cdot\|_{w}$-projection onto the space $A$, and $U$ abbreviates $U(\theta_{0};Z)$. The second term on the right of (10) is non-negative, establishing $V_{1}\leq V_{3}$.
@@
$V_{3}\;$ $=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}$
$\qquad\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\left\{U(\theta_{0};Z)-b(O)\right\}^{\otimes 2}\right]$
@@
Corollary 3 identifies the main advantage of the calibration approach over classical AIPW. Standard AIPW augments the estimating equation using only the working outcome-regression component $b(O)$. By contrast, the generalized entropy calibration estimator projects onto the larger space generated by both $b(O)$ and the debiasing covariates induced by the propensity-score model. When the OR model is misspecified, this enlarged projection space can substantially reduce the residual variance, yielding a strict efficiency gain without changing the basic doubly robust logic. This variance-reduction property distinguishes GEC from both TMLE (van der Laan and Rose, 2011) and the regularized calibrated estimation of Tan (2020).
@@
$V_{1}^{*}\;$ $=\;\mathrm{Var}\bigl\{U(\theta_{0};Z)\bigr\}$
$\qquad\;+\;\mathbb{E}\!\left[\left(\frac{1}{\pi(O;\varphi_{0})}-1\right)\mathrm{Var}\{U(\theta_{0};Z)\mid O\}\right],$
@@
5 Computation and Cross-Fitting
The proposed framework is defined through calibration constraints, but its practical implementation is straightforward. The main computational feature distinguishing the present setting from standard entropy calibration is that the balancing constraint depends on the target parameter $\theta$. As a result, the calibration weights themselves depend on $\theta$, and estimation proceeds through a profile optimization: for a candidate value of $\theta$, we first solve the calibration problem to obtain the corresponding weights, and we then update $\theta$ using the resulting weighted estimating equation. Thus, the method combines a low-dimensional convex calibration step with an outer parameter-update step.
This nested structure is the computational counterpart of the statistical formulation in Section 3. The inner step enforces the balancing and debiasing constraints through generalized entropy calibration, while the outer step updates the parameter of interest using the calibrated weighted estimating equation. Because the inner problem is convex in the dual variable $\lambda$, each calibration update is computationally stable and can be solved with standard optimization routines.
Algorithm 1 summarizes the resulting two-loop profile optimization procedure for computing $\hat{\theta}_{\omega}$.
@@ 5.1 Profile optimization
$\hat{\theta}_{\omega}=\arg\max_{\theta}\min_{\lambda}\;\rho_{G}(\lambda,\theta).$ (12)
Equivalently, defining the profile objective $L(\theta)=\rho_{G}(\hat{\lambda}(\theta),\theta)$ where $\hat{\lambda}(\theta)=\arg\min_{\lambda}\rho_{G}(\lambda,\theta)$, we have
$\hat{\theta}_{\omega}=\arg\max_{\theta}L(\theta).$
We solve (12) using a two-loop procedure. For a fixed $\theta$, the inner loop minimizes $\rho_{G}(\lambda,\theta)$ over $\lambda$ using Newton’s method. The outer loop updates $\theta$ via a Newton or BFGS step based on the gradient of $L(\theta)$. By the envelope theorem (Danskin, 2012), $\nabla_{\theta}L(\theta)=\nabla_{\theta}\rho_{G}(\hat{\lambda}(\theta),\theta)$, since $\nabla_{\lambda}\rho_{G}=0$ at the inner optimum. The procedure is summarized in Algorithm 1.
Algorithm 1 Two-Loop Profile Optimization for $\hat{\theta}_{\omega}$. At a high level, the algorithm alternates between calibrating the weights for the current value of $\theta$ and updating $\theta$ using the resulting profile objective.
@@
5:  Outer loop: Update
$\theta^{(k+1)}\leftarrow\theta^{(k)}-[\nabla^{2}_{\theta}L(\theta^{(k)})]^{-1}\,\nabla_{\theta}L(\theta^{(k)})$
@@ 5.2 Calibration via cross-fitting
In practice, the ideal calibration function
$b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$
is unknown and must be approximated from the data. This is the point at which modern prediction enters the framework. Rather than specifying a parametric outcome-regression model, we approximate $b^{*}(\theta;O_{i})$ by predicting the missing component $M_{i}$ from the observed component $O_{i}$ and then plugging that prediction into the estimating function:
$b^{*}(\theta;O_{i})\approx U(\theta;O_{i},\hat{M}_{i}).$
This yields a prediction-powered version of generalized entropy calibration in which the balancing constraint is constructed from data-adaptive imputations of the missing part of the full-data vector.
To avoid the overfitting bias that would arise if the same observations were used both to train the prediction rule and to define the calibration function, we adopt $K$-fold cross-fitting, following Chernozhukov et al. (2018) and Angelopoulos et al. (2023). Cross-fitting ensures that each prediction $\hat{M}_{i}$ is obtained from a model trained on data that exclude unit $i$, thereby preserving the separation between the learning step and the estimating step.
Let $\mathcal{I}=\{1,\ldots,N\}$ and $\mathcal{S}=\{i\in\mathcal{I}:\delta_{i}=1\}$. We randomly partition $\mathcal{I}$ into $K$ disjoint folds $\{\mathcal{I}^{(k)}\}_{k=1}^{K}$. For each fold $k$, we fit a prediction model $\hat{m}^{(-k)}(\cdot)$ using training data $\mathcal{S}^{(-k)}=\mathcal{S}\setminus(\mathcal{S}\cap\mathcal{I}^{(k)})$ and compute out-of-fold predictions $\hat{M}_{i}^{(-k)}=\hat{m}^{(-k)}(O_{i})$ for $i\in\mathcal{I}^{(k)}$.
Using these out-of-fold predictions, we evaluate the calibration function for every unit in the sample, including units with $\delta_{i}=0$. This is essential because the balancing constraint is imposed at the full-sample level: although the weights are attached only to observed units, the calibration targets themselves must be computed for all units.
Using these cross-fitted predictions, we solve the calibration problem:
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (13)
$\text{subject to}\quad\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}\delta_{i}\,\omega_{i}\,U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right)$
$\hskip 99.58464pt=\sum_{k=1}^{K}\sum_{i\in\mathcal{I}^{(k)}}U\!\left(\theta;\,O_{i},\hat{M}_{i}^{(-k)}\right),$ (14)
@@
This cross-fitted formulation preserves the same two-part logic as the original calibration problem. The balancing constraint uses out-of-fold predictions to approximate the conditional mean structure of the estimating function, while the debiasing constraint continues to anchor the weights to the propensity-score model. Thus, cross-fitting does not alter the inferential architecture of the method; it simply replaces a parametric calibration function by a flexible learned approximation.
Note that the prediction model $\hat{m}^{(-k)}$ is trained only on the complete cases in $\mathcal{S}^{(-k)}$, but the out-of-fold predictions $\hat{M}_{i}^{(-k)}$ are computed for all units $i\in\mathcal{I}^{(k)}$, including those with $\delta_{i}=0$. This is essential: the balancing constraint (14) involves the full sample, so calibration function values are needed for every unit regardless of whether $M_{i}$ is observed.
@@
Under correct specification of the propensity-score model, the cross-fitted predictor $\hat{M}^{(-k)}$ needs only to consistently approximate the conditional mean of the missing component in mean square:
$E\|\hat{M}_{i}^{(-k)}-E(M_{i}\mid O_{i})\|^{2}=o(1).$
No specific rate such as $N^{-1/4}$ is required for first-order validity along this PS-correct pathway, because the debiasing constraint anchors the calibration weights to the IPW solution even when the predictor converges relatively slowly. Stronger conditions may be needed when first-order validity relies instead on the OR pathway; see Lee and Kim (2026) for more technical details.